gross national product birr 5000, indirect taxes birr 580, net domestic product birr 5900, net national product birr 4250 the a) calculate gross domestic product b) net national income at factor cost c) calculate net factor income


1
Expert's answer
2022-02-10T08:24:29-0500

Qn. A

To calculate the value of GDP, we know the following;

"NNP=4250"

"GNP=5000"

"NDP=5900"

"NNP=GNP-depreciation"

"4250=5000-depreciation"

"depreciation=5000-4250=750"

"NDP=GDP-depreciation"

"5900=GDP-750"

therefore;

"GDP=5900+750"

"\bold{GDP=6650}"


Qn. B

Net national income at factor is given as;

"NNI=GDP+NX-depreciation"

"NNI=6650+NX-750"

"NNI=5900+NX"

"NX =net\ factor\ income"

but we know;

"GNP=GDP+NX"

"5000=6650+NX"

"NX=-1650"

"NNI=5900+(-1650)"

"\bold{NNI=4250}"


Qn. C

Net factor income "=NX"

from our calculations above, "NX=-1650"

"\therefore \bold{Net\ factor\ income =-1650}"